# Credal Prediction based on Relative Likelihood

### v1

@@ Appendix A Extended Related Work
Uncertainty Representation. In machine learning, uncertainty is often represented by Bayesian methods [Mackay, 1992, Blundell et al., 2015]. Frequently, the Bayesian posterior is approximated by ensembles [Lakshminarayanan et al., 2017] or methods such as Dropout [Gal and Ghahramani, 2016] or Laplace approximation [Daxberger et al., 2021]. An important characteristic of such representations, especially for uncertainty tasks, is diversity [D’Angelo and Fortuin, 2021, Wood et al., 2023]. Some works enforce this by means of regularization [de Mathelin et al., 2023], whereas others vary hyperparameters across models within the ensemble [Wenzel et al., 2020] or enforce diversity in the representations [Lopes et al., 2022].
@@
Alternatively, credal sets have been used in the fields of imprecise probability and machine learning to represent model uncertainty [Zaffalon, 2001, Corani and Zaffalon, 2008, Corani and Mignatti, 2015]. Antonucci et al. [2012a] proposed to generate such sets based on relative likelihoods. Relative likelihoods, also referred to as normalized likelihoods, have also been used in machine learning with simple model classes such as logistic regression [Senge et al., 2014, Cella and Martin, 2024]. In this work, we build upon these approaches and address the challenges that emerge when adapting the relative likelihood to a setting with complex predictors.
@@
Recently, credal sets have been applied in the context of machine learning. Wang et al. [2024b] take multiple samples from a Bayesian posterior or ensembles and derive class-wise lower and upper probabilities from this. Based on these samples or ensemble predictions, a credal set is constructed by including all probability distributions such that the individual class probabilities are in the respective lower to upper class probability interval. Nguyen et al. [2025] also use ensemble predictions, but they additionally allow for the discarding of potential outliers to ensure that the resulting credal set too large. This is done by comparing all ensemble predictions to a representative prediction, e.g. the mean prediction, using some distance between distributions, and keeping only $(1-\alpha)\cdot 100\%$ closest predictions. The credal set is then constructed by taking the convex hull of the remaining probability distributions. Other methods directly train neural networks to predict intervals by explicitly having the neural network predict a lower and upper probability for every class [Wang et al., 2024a]. In combination with a custom loss function, consisting of regular cross-entropy for the upper probabilities and cross-entropy computed on the highest loss subset of a batch for the lower probabilities, the claim is that this approach encourages both “optimistic” and “pessimistic” predictions. Credal sets are constructed by taking the same interval-based approach as [Wang et al., 2024b]. Besides this, hybrid methods, combining multiple uncertainty frameworks, have also been proposed. Caprio et al. [2023] combine Bayesian deep learning and credal sets by considering sets of priors over weights of neural networks. Training these neural network by variational inference then results in a set of posteriors. Based on this, a credal set is constructed by sampling from the Bayesian neural networks and taking the convex hull of the sampled probability distributions. Another approach leverages conformal prediction to construct credal sets with validity guarantees [Javanmardi et al., 2024]. However, this work uses distributions over classes to perform the conformal calibration step. Since our method does not require such data, we do not use this method as a baseline. Our work distinguishes itself from aforementioned works by making use of relative likelihood cuts which allow for an intuitive and adaptable construction of the credal set.
@@
Uncertainty Quantification. Given a credal representation of uncertainty, there are many ways to quantify uncertainty. Some measures consider only the epistemic uncertainty [Abellán and Moral, 2000], whereas others use entropy to reason about the total uncertainty [Abellán and Moral, 2003]. In this line of work, Abellán et al. [2006] also proposed measures based on entropy that decompose the total uncertainty into an aleatoric and epistemic component. Conversely, Antonucci et al. [2012b] quantify uncertainty by measuring the lack of dominance of one class over others in the predictive distributions represented by the credal set. Hüllermeier et al. [2022] offer a critical analysis of these measures and propose an alternative for the dominance-based measure. Recently, Hofman et al. [2024] proposed to quantify credal uncertainty based on a decomposition of scoring rules. In the following, we consider the uncertainty measures by Abellán et al. [2006] in order to ensure a fair comparison with previous works.
@@ B.1 Datasets
ChaosNLI The ChaosNLI dataset, introduced by Nie et al. [2020], is a large-scale dataset designed to study human disagreement in natural language inference (NLI) tasks. It comprises 100 human annotations per example for 3,113 examples from the SNLI and MNLI datasets, and 1,532 examples from the $\alpha$NLI dataset, totaling approximately 464,500 annotations. In line with Javanmardi et al. [2024], we use only the SNLI and MNLI subsets of the ChaosNLI dataset, but for simplicity, we will refer to this as the ChaosNLI dataset. Each example includes metadata such as the unique identifier, counts of each label assigned by annotators, the majority label, label distribution, entropy of the label distribution, the original example text, and the original label from the source dataset. This dataset enables a detailed analysis of the distribution of human opinions in NLI tasks, highlighting instances of high disagreement and questioning the validity of using majority labels as the sole ground truth. ChaosNLI is publicly available under the Creative Commons Attribution-NonCommercial 4.0 International (CC BY-NC 4.0) license. We train our models on the 768-dimensional embeddings of the ChaosNLI dataset retrieved from https://github.com/alireza-javanmardi/conformal-credal-sets. We refer to [Javanmardi et al., 2024] for more details on the generation of the embeddings.
@@
CIFAR-10 The CIFAR-10 dataset is a widely used benchmark in machine learning and computer vision, introduced by Krizhevsky et al. [2009], and Geoffrey Hinton in 2009. It comprises 60,000 color images at a resolution of 32×32 pixels, evenly distributed across 10 distinct classes: airplane, automobile, bird, cat, deer, dog, frog, horse, ship, and truck. The dataset is partitioned into 50,000 training images and 10,000 test images, organized into five training batches and one test batch, each containing 10,000 images. The dataset is publicly available and has been utilized extensively for developing and benchmarking machine learning models. While the original dataset does not specify a license, various distributions, such as those provided by TensorFlow Datasets, are released under the Creative Commons Attribution 4.0 License.
@@
CIFAR-10H The CIFAR-10H dataset provides human-derived soft labels for the 10,000 images in the CIFAR-10 test set, capturing the variability in human annotation during image classification tasks. Developed by Peterson et al. [2019], the dataset comprises 511,400 annotations collected from 2,571 Amazon Mechanical Turk workers, with each image receiving approximately 51 labels. Annotators classified images into one of the ten CIFAR-10 categories, enabling the construction of probability distributions over labels for each image. CIFAR-10H is publicly available under the Creative Commons BY-NC-SA 4.0 license.
@@
CIFAR-100 The CIFAR-100 dataset, introduced by Krizhevsky et al. [2009], comprises 60,000 color images at 32×32 resolution, divided into 100 classes with 600 images each. Each image has a “fine” label (specific class) and a “coarse” label (superclass), with the 100 classes grouped into 20 superclasses. The dataset is split into 50,000 training and 10,000 test images. It is a subset of the Tiny Images dataset and is commonly used for evaluating image classification algorithms. While the original dataset does not specify a license, various distributions, such as those provided by TensorFlow Datasets, are released under the Creative Commons Attribution 4.0 License.
@@
QualityMRI The QualityMRI dataset, introduced by Obuchowicz et al. [2020], is part of the Data-Centric Image Classification (DCIC) Benchmark, which aims to evaluate the impact of dataset curation on model performance. The dataset contains 310 magnetic resonance (MRI) images spanning various quality levels and is designed to assess the MRI image quality. The dataset is publicly available under the Creative Commons BY-SA 4.0 license.
@@
SVHN The SVHN dataset, introduced by Netzer et al. [2011], consists of over 600,000 32×32 RGB images of digits (0–9) obtained from real-world house number images in Google Street View. It includes three subsets: 73,257 training images, 26,032 test images, and 531,131 additional images for extra training. The dataset is designed for digit recognition tasks with minimal preprocessing. While the original dataset does not specify a license, various distributions, such as those provided by TensorFlow Datasets, are released under the Creative Commons Attribution 4.0 License.
@@
Places365 Places365, introduced by Zhou et al. [2018], is a large-scale scene recognition dataset containing 1.8 million training images across 365 scene categories. The validation set includes 50 images per category, and the test set has 900 images per category. An extended version, Places365-Challenge-2016, adds 6.2 million images and 69 new scene classes, totaling 8 million images over 434 categories. While the original dataset does not specify a license, various distributions, such as those provided by TensorFlow Datasets, are released under the Creative Commons Attribution 4.0 License.
@@
FMNIST Fashion-MNIST (FMNIST), introduced by Xiao et al. [2017], is a dataset of Zalando’s article images, comprising 70,000 28×28 grayscale images labeled across 10 classes, such as T-shirt/top, Trouser, and Sneaker. It includes 60,000 training and 10,000 test images and serves as a direct replacement for the original MNIST dataset for benchmarking machine learning algorithms. FMNIST is publicly available under the MIT License.
@@
ImageNet ImageNet, introduced by Deng et al. [2009], is a large-scale image database organized according to the WordNet hierarchy, containing over 14 million images across more than 20,000 categories. The ILSVRC subset (ImageNet-1K) includes 1,281,167 training images, 50,000 validation images, and 100,000 test images across 1,000 classes. The dataset is available for free to researchers for non-commercial use.
@@ B.2 Models
Fully-Connected Network We train fully-connected neural networks on the ChaosNLI dataset. The network consists of 4 linear layers with $[768-256-64-16-3]$ units with ReLU activations, except for the last layer, which has Softmax transformation to transform the logits into probabilities. We use the hyperparameters (cf. Table 2) similar to the optimal parameters found in Javanmardi et al. [2024].
ResNet18 For experiments on the CIFAR-10 dataset, we use the PyTorch ResNet-18 implementation and hyperparameters provided by https://github.com/kuangliu/pytorch-cifar. This model is specifically optimized for CIFAR-10 and is trained from scratch, without any pretraining on ImageNet. For experiments on the QualityMRI dataset, we use the ResNet18 implementation from the PyTorch torchvision package with random initialization, i.e. no pretrained weights.
Hyperparameters Each dataset is trained using a dedicated set of hyperparameters as presented in Table 2. We evaluated multiple configurations and selected the best-performing ones for each dataset. To ensure fair and consistent comparisons, all models trained on a given dataset, both our approach and the baselines, use the same hyperparameter settings. The only exception is the CreBNN, which requires a KL-divergence penalty of $71e-$ 7 and zero weight decay when using the Adam optimizer [Kingma and Ba, 2015]. When we apply the SGD optimizer with a learning rate scheduler, namely Cosine Annealing [Loshchilov and Hutter, 2017], CreBNN requires additionally a momentum of $0.9$ to enable effective learning.
@@ B.3 Out-of-Distribution Detection
We use the SVHN [Netzer et al., 2011], Places365 [Zhou et al., 2018], CIFAR-100 [Krizhevsky et al., 2009], Fashion-MNIST [Xiao et al., 2017], and ImageNet [Le and Yang, 2015] datasets for Out-of-Distribution detection.
@@ B.4 Computing Uncertainty
Computing the total and aleatoric uncertainty involves solving a per instance optimization problem, after which the epistemic uncertainty is obtained as their difference (see Equation 5). This optimization is performed using SciPy’s minimize function with the SLSQP solver and the default parameters.
Interval-Based For the interval-based credal sets, the optimization is initialized with the mean of the predicted distributions, bounded between the lower and upper probabilities for each class, and constrained to ensure that the solution forms a valid probability distribution, namely, the class probabilities must sum to $1$.
Convex Hull For the credal sets based on the convex hull, the optimization is done on the weights of convex combination, instead of the probability distribution. Uniform weights are used as the initial value, the weights are bounded between 00 and $1$, and constrained to sum to $1$.
Estimated Computing Time Here, we provide an estimated upper bound of the computation time for computing the lower and upper entropy of a credal set based on the example of the largest possible credal set, namely the full probability simplex for a $10$-class problem, such as CIFAR-10. For this largest set, computing the upper entropy using an interval-based credal set takes on average $0.03$ seconds, while computing the lower entropy takes about $0.02$ seconds. For the convex hull-based credal set of the same size, the average computation time is $0.07$ seconds for upper entropy and $0.03$ seconds for lower entropy. To simplify our estimation, we average these times and assume that for a single instance it takes $0.03$ seconds to optimize the lower entropy and $0.04$ seconds to optimize upper entropy, summing up to $0.07$ seconds of computing time per instance.
@@ B.6 Baselines
Credal Wrapper (CreWra) The Credal Wrapper was initially implemented in TensorFlow, but we reimplemented it in PyTorch to ensure compatibility with our framework. It follows a standard ensemble learning approach, training multiple models independently. Like our method, the Credal Wrapper constructs credal sets using class-wise upper and lower probability bounds, making it well-aligned with our implementation. Overall, we closely follow both the original paper and their available implementation [Wang et al., 2024b].
Credal Ensembling ($\text{CreEns}_{\alpha}$) Since no official code was available for neural network implementations of Credal Ensembling, we reimplemented the method ourselves. Our implementation closely follows all details provided in Nguyen et al. [2025]. The approach builds on standard ensemble training procedures, with inference adapted according to their proposed method of sorting predictions based on a distance measure and selecting only $\alpha\%$ closest predictions to construct credal sets. In our experiments, we use the Euclidean distance measure and evaluate several values of $\alpha$.
Credal Deep Ensembles (CreNet) As the official implementation of Credal Deep Ensembles is only available in TensorFlow, we reimplemented the method in PyTorch to ensure compatibility with our codebase. Our implementation closely mirrors the original TensorFlow code, particularly in adapting the model architecture and loss function. Specifically, we replace each model’s final linear layer with a head comprising a linear layer outputting $2\times\text{classes}$ (representing upper and lower probability bounds), followed by a batch normalization layer and the custom IntSoftmax layer. We also reimplemented the proposed loss function, which computes a cross-entropy loss for the upper bounds and selectively backpropagates the lower-bound loss only for the $\delta\%$ of samples with the highest loss values, as described in Wang et al. [2024a]. In our experiments, we use $\delta=0.5$, as suggested in Wang et al. [2024a].
Credal Bayesian Deep Learning (CreBNN) No code or implementation details for Credal Bayesian Deep Learning (CreBNN) were made publicly available, and despite multiple attempts to contact the authors, we received no response or further clarification. As a result, we reimplemented the method ourselves based solely on the high-level description provided in the paper. In our implementation, each ensemble member is a Bayesian neural network (BNN) trained with variational inference using different priors, with prior means $\mu$ sampled from $[-1,1]$ and standard deviations $\sigma$ from $[0.1,2]$ to form a diverse prior set. During inference, we draw one sample from each BNN to obtain a finite set of probability distributions, and construct the credal set as the convex hull of these predictions.

### v2

@@ Appendix A Extended Related Work
Uncertainty Representation. In machine learning, uncertainty is often represented by Bayesian methods [mackayBayesianMethods1992, blundellWeightUncertainty2015]. Frequently, the Bayesian posterior is approximated by ensembles [lakshminarayananDeepEnsembles2017] or methods such as Dropout [galDropoutAs2016] or Laplace approximation [daxbergerLaplaceRedux2021]. An important characteristic of such representations, especially for uncertainty tasks, is diversity [d2021repulsive, woodUnifiedTheory2023]. Some works enforce this by means of regularization [deMathelinDeepAnti2023], whereas others vary hyperparameters across models within the ensemble [wenzelHyperparameterEnsembles2020] or enforce diversity in the representations [lopesNoOne2022].
@@
Alternatively, credal sets have been used in the fields of imprecise probability and machine learning to represent model uncertainty [zaffalonStatisticalInference2001, coraniLearningReliable2008, coraniCredalModel2015]. antonucciLikelihoodBased2012 proposed to generate such sets based on relative likelihoods. Relative likelihoods, also referred to as normalized likelihoods, have also been used in machine learning with simple model classes such as logistic regression [sengeReliableClassification2014, cellaVariationalApproximations2024]. In this work, we build upon these approaches and address the challenges that emerge when adapting the relative likelihood to a setting with complex predictors.
@@
Recently, credal sets have been applied in the context of machine learning. wangCredalWrapper2024 take multiple samples from a Bayesian posterior or ensembles and derive class-wise lower and upper probabilities. Based on these samples, a credal set is constructed by including all probability distributions such that the individual predicted class probabilities are in the respective lower to upper class probability interval. nguyenCredalEnsembling2025 also construct credal sets from ensemble predictions, but with the additional option of discarding potential outliers to prevent the set from becoming too large. This is done by comparing all ensemble predictions to a representative prediction, e.g. the mean prediction, using some distance between distributions, and keeping only $(1-\alpha)\cdot 100\%$ closest predictions. The credal set is then constructed by taking the convex hull of the remaining probability distributions. Other methods directly train neural networks to predict intervals by explicitly predicting a lower and upper probability for every class [wangCredalDeep2024]. In combination with a custom loss function, consisting of regular cross-entropy for the upper probabilities and cross-entropy computed on the highest loss subset of a batch for the lower probabilities, the claim is that this approach encourages both “optimistic” and “pessimistic” predictions. Credal sets are constructed by taking the same interval-based approach as [wangCredalWrapper2024]. Besides this, hybrid methods, combining multiple uncertainty frameworks, have also been proposed. caprioImpreciseBayesian2023 combine Bayesian deep learning and credal sets by considering sets of priors over weights of neural networks. Training these neural network by variational inference then results in a set of posteriors. Based on this, a credal set is constructed by sampling from the Bayesian neural networks and taking the convex hull of the sampled probability distributions. Another approach leverages conformal prediction to construct credal sets with validity guarantees [javanmardiConformalizedCredal2024]. However, this work uses distributions over classes to perform the conformal calibration step. Since our method does not require such data, we exclude this method as a baseline. Our work distinguishes itself from aforementioned works by using relative likelihood cuts which allow for an intuitive and adaptable construction of the credal set.
Moreover, our approach is conceptually related to Rashomon sets [semenova2022existence]. Both characterize a collection of models whose performance exceeds a given threshold, thereby facing the same challenge in approximating this set of plausible models [donnelly2025rashomon]. However, the objectives differ: Rashomon sets are primarily concerned with interpretability and (syntactic) model diversity, whereas our focus is on uncertainty quantification and predictive diversity.
@@
Uncertainty Quantification. Given a credal representation of uncertainty, there are many ways to quantify uncertainty. Some measures consider only the epistemic uncertainty [abellanNonSpecificty2000], whereas others use entropy to reason about the total uncertainty [abellanMaximumOf2003]. In this line of work, abellanDisaggregatedTotal2006 also proposed measures based on entropy that decompose the total uncertainty into an aleatoric and epistemic component. Conversely, antonucciActiveLearning2012 quantify uncertainty by measuring the lack of dominance of one class over others in the predictive distributions represented by the credal set. huellermeierQuantificationOf2022 offer a critical analysis of these measures and propose an alternative for the dominance-based measure. Recently, hofmanQuantifyingAleatoric2024 proposed to quantify credal uncertainty based on a decomposition of scoring rules. In the following, we consider the uncertainty measures by abellanDisaggregatedTotal2006 in order to ensure a fair comparison with previous works.
@@ B.1 Datasets
ChaosNLI
[…]
CIFAR-10
@@
The CIFAR-10 dataset is a widely used benchmark in machine learning and computer vision, introduced by krizhevsky2009learning, and Geoffrey Hinton in 2009. It comprises 60,000 color images at a resolution of 32×32 pixels, evenly distributed across 10 distinct classes: airplane, automobile, bird, cat, deer, dog, frog, horse, ship, and truck. The dataset is partitioned into 50,000 training images and 10,000 test images, organized into five training batches and one test batch, each containing 10,000 images. The dataset is publicly available and has been utilized extensively for developing and benchmarking machine learning models. While the original dataset does not specify a license, various distributions, such as those provided by TensorFlow Datasets, are released under the Creative Commons Attribution 4.0 License.
CIFAR-10H
@@
The CIFAR-10H dataset provides human-derived soft labels for the 10,000 images in the CIFAR-10 test set, capturing the variability in human annotation during image classification tasks. Developed by petersonHumanUncertainty2019, the dataset comprises 511,400 annotations collected from 2,571 Amazon Mechanical Turk workers, with each image receiving approximately 51 labels. Annotators classified images into one of the ten CIFAR-10 categories, enabling the construction of probability distributions over labels for each image. CIFAR-10H is publicly available under the Creative Commons BY-NC-SA 4.0 license.
CIFAR-100
@@
The CIFAR-100 dataset, introduced by krizhevsky2009learning, comprises 60,000 color images at 32×32 resolution, divided into 100 classes with 600 images each. Each image has a “fine” label (specific class) and a “coarse” label (superclass), with the 100 classes grouped into 20 superclasses. The dataset is split into 50,000 training and 10,000 test images. It is a subset of the Tiny Images dataset and is commonly used for evaluating image classification algorithms. While the original dataset does not specify a license, various distributions, such as those provided by TensorFlow Datasets, are released under the Creative Commons Attribution 4.0 License.
QualityMRI
@@
The QualityMRI dataset, introduced by qualityMRI, is part of the Data-Centric Image Classification (DCIC) Benchmark, which aims to evaluate the impact of dataset curation on model performance. The dataset contains 310 magnetic resonance (MRI) images spanning various quality levels and is designed to assess the MRI image quality. The dataset is publicly available under the Creative Commons BY-SA 4.0 license.
SVHN
@@
The SVHN dataset, introduced by netzerReading2011, consists of over 600,000 32×32 RGB images of digits (0–9) obtained from real-world house number images in Google Street View. It includes three subsets: 73,257 training images, 26,032 test images, and 531,131 additional images for extra training. The dataset is designed for digit recognition tasks with minimal preprocessing. While the original dataset does not specify a license, various distributions, such as those provided by TensorFlow Datasets, are released under the Creative Commons Attribution 4.0 License.
Places365
@@
Places365, introduced by zhouPlaces2018, is a large-scale scene recognition dataset containing 1.8 million training images across 365 scene categories. The validation set includes 50 images per category, and the test set has 900 images per category. An extended version, Places365-Challenge-2016, adds 6.2 million images and 69 new scene classes, totaling 8 million images over 434 categories. While the original dataset does not specify a license, various distributions, such as those provided by TensorFlow Datasets, are released under the Creative Commons Attribution 4.0 License.
FMNIST
@@
Fashion-MNIST (FMNIST), introduced by xiaoFashionMNIST2017, is a dataset of Zalando’s article images, comprising 70,000 28×28 grayscale images labeled across 10 classes, such as T-shirt/top, Trouser, and Sneaker. It includes 60,000 training and 10,000 test images and serves as a direct replacement for the original MNIST dataset for benchmarking machine learning algorithms. FMNIST is publicly available under the MIT License.
ImageNet
@@
ImageNet, introduced by deng2009imagenet, is a large-scale image database organized according to the WordNet hierarchy, containing over 14 million images across more than 20,000 categories. The ILSVRC subset (ImageNet-1K) includes 1,281,167 training images, 50,000 validation images, and 100,000 test images across 1,000 classes. The dataset is available for free to researchers for non-commercial use.
@@ B.2 Models
Fully-Connected Network
We train fully-connected neural networks on the ChaosNLI dataset. The network consists of 4 linear layers with [76825664163][768-256-64-16-3] units with ReLU activations, except for the last layer, which has Softmax transformation to transform the logits into probabilities. We use the hyperparameters (cf. Table˜2) similar to the optimal parameters found in javanmardiConformalizedCredal2024.
ResNet18
For experiments on the CIFAR-10 dataset, we use the PyTorch ResNet-18 implementation and hyperparameters provided by https://github.com/kuangliu/pytorch-cifar. This model is specifically optimized for CIFAR-10 and is trained from scratch, without any pretraining on ImageNet. For experiments on the QualityMRI dataset, we use the ResNet18 implementation from the PyTorch torchvision package with random initialization, i.e. no pretrained weights.
Hyperparameters
Each dataset is trained using a dedicated set of hyperparameters as presented in Table˜2. We evaluated multiple configurations and selected the best-performing ones for each dataset. To ensure fair and consistent comparisons, all models trained on a given dataset, both our approach and the baselines, use the same hyperparameter settings. The only exception is the CreBNN, which requires a KL-divergence penalty of 1e71e-7 and zero weight decay when using the Adam optimizer [kingmaAdam2015]. When we apply the SGD optimizer with a learning rate scheduler, namely Cosine Annealing [loshchilovSGDR2017], CreBNN requires additionally a momentum of $0.9$ to enable effective learning.
@@ B.3 Out-of-Distribution Detection
We use the SVHN [netzerReading2011], Places365 [zhouPlaces2018], CIFAR-100 [krizhevsky2009learning], Fashion-MNIST [xiaoFashionMNIST2017], and ImageNet [leTinyImagenet2015] datasets for Out-of-Distribution detection.
@@ B.4 Computing Uncertainty
Computing the total and aleatoric uncertainty involves solving a per instance optimization problem, after which the epistemic uncertainty is obtained as their difference (see Equation˜5). This optimization is performed using SciPy’s minimize function with the SLSQP solver and the default parameters.
Interval-Based
For the interval-based credal sets, the optimization is initialized with the mean of the predicted distributions, bounded between the lower and upper probabilities for each class, and constrained to ensure that the solution forms a valid probability distribution, namely, the class probabilities must sum to $1$.
Convex Hull
For the credal sets based on the convex hull, the optimization is done on the weights of convex combination, instead of the probability distribution. Uniform weights are used as the initial value, the weights are bounded between 0 and $1$, and constrained to sum to $1$.
Estimated Computing Time
Here, we provide an estimated upper bound of the computation time for computing the lower and upper entropy of a credal set based on the example of the largest possible credal set, namely the full probability simplex for a $10$-class problem, such as CIFAR-10. For this largest set, computing the upper entropy using an interval-based credal set takes on average $0.03$ seconds, while computing the lower entropy takes about $0.02$ seconds. For the convex hull-based credal set of the same size, the average computation time is $0.07$ seconds for upper entropy and $0.03$ seconds for lower entropy. To simplify our estimation, we average these times and assume that for a single instance it takes $0.03$ seconds to optimize the lower entropy and $0.04$ seconds to optimize upper entropy, summing up to $0.07$ seconds of computing time per instance.
@@ B.6 Baselines
Credal Wrapper (CreWra)
The Credal Wrapper was initially implemented in TensorFlow, but we reimplemented it in PyTorch to ensure compatibility with our framework. It follows a standard ensemble learning approach, training multiple models independently. Like our method, the Credal Wrapper constructs credal sets using class-wise upper and lower probability bounds, making it well-aligned with our implementation. Overall, we closely follow both the original paper and their available implementation [wangCredalWrapper2024].
Credal Ensembling ($\text{CreEns}_{\alpha}$)
Since no official code was available for neural network implementations of Credal Ensembling, we reimplemented the method ourselves. Our implementation closely follows all details provided in nguyenCredalEnsembling2025. The approach builds on standard ensemble training procedures, with inference adapted according to their proposed method of sorting predictions based on a distance measure and selecting only $\alpha\%$ closest predictions to construct credal sets. In our experiments, we use the Euclidean distance measure and evaluate several values of $\alpha$.
Credal Deep Ensembles (CreNet)
As the official implementation of Credal Deep Ensembles is only available in TensorFlow, we reimplemented the method in PyTorch to ensure compatibility with our codebase. Our implementation closely mirrors the original TensorFlow code, particularly in adapting the model architecture and loss function. Specifically, we replace each model’s final linear layer with a head comprising a linear layer outputting $2\times\text{classes}$ (representing upper and lower probability bounds), followed by a batch normalization layer and the custom IntSoftmax layer. We also reimplemented the proposed loss function, which computes a cross-entropy loss for the upper bounds and selectively backpropagates the lower-bound loss only for the $\delta\%$ of samples with the highest loss values, as described in wangCredalDeep2024. In our experiments, we use $\delta=0.5$, as suggested in wangCredalDeep2024.
Credal Bayesian Deep Learning (CreBNN)
No code or implementation details for Credal Bayesian Deep Learning (CreBNN) were made publicly available, and despite multiple attempts to contact the authors, we received no response or further clarification. As a result, we reimplemented the method ourselves based solely on the high-level description provided in the paper. In our implementation, each ensemble member is a Bayesian neural network (BNN) trained with variational inference using different priors, with prior means $\mu$ sampled from [1,1][-1,1] and standard deviations $\sigma$ from $[0.1,2]$ to form a diverse prior set. During inference, we draw one sample from each BNN to obtain a finite set of probability distributions, and construct the credal set as the convex hull of these predictions.